# Sphere Decoding Revisited

### v1

@@ Abstract
In this paper, the paradigm of sphere decoding (SD) for solving the integer least square problem (ILS) is revisited, where extra degrees of freedom are introduced to exploit the decoding potential. Firstly, the equivalent sphere decoding (ESD) is proposed, which is essentially the same with the classic Fincke-Pohst sphere decoding but characterizes the sphere radius $D>0$ with two new parameters named as initial searching size $K>1$ and deviation factor $\sigma>0$. By fixing $\sigma$ properly, we show that given the sphere radius $D\triangleq\sigma\sqrt{2\ln K}$, the complexity of ESD in terms of the number of visited nodes is upper bounded by $|S|<nK$, thus resulting in an explicit and tractable decoding trade-off solely controlled by $K$. To the best of our knowledge, this is the first time that the complexity of sphere decoding is exactly specified, where considerable decoding potential can be explored from it. After that, two enhancement mechanisms named as normalized weighting and candidate protection are proposed to further upgrade the ESD algorithm. On one hand, given the same setups of $K$ and $\sigma$, a larger sphere radius is achieved, indicating a better decoding trade-off. On the other hand, the proposed ESD algorithm is generalized, which bridges suboptimal and optimal decoding performance through the flexible choice of $K$. Finally, further performance optimization and complexity reduction with respect to ESD are also derived, and the introduced tractable and flexible decoding trade-off is verified through large-scale MIMO detection.
@@ I Introduction
Nowadays, the large-scale multiple-input multiple-output (MIMO) system has become a promising extension of MIMO in 5G, which boosts the network capacity on a much greater scale without extra bandwidth [5595728, 6736761, 6375940, 6457363]. However, the dramatically increased system size also places a pressing challenge on signal detection, which belongs to integer least square (ILS) problem. As a general way to realize the maximum-likelihood (ML) decoding, the traditional sphere decoding (SD) turns out to be impractical due to the unaffordable complexity in large-scale systems [Kannan, Agrell2002, HassibiExpected, DamenDetectionSearch]. As for those near-ML decoding schemes like fixed-complexity sphere decoding (FCSD), K-best decoder, etc., they are also inapplicable due to the intensive complexity increment and terrible performance deterioration [ZhengWangTSP2019, Hassibisoft, JaldenFSD, 4154776, EmbeddingLuzzi]. To this end, most of related works try to focus on the low-complexity decoding schemes although their performance are severely limited [TabuSrinidhi, DaiL1, GaoX1, MSMIMO1, 6008619, ZhengWangTCOM2019].
@@
First of all, equivalent sphere decoding (ESD) is proposed to solve the ILS problem (i.e., CVP in lattice decoding) in large-scale MIMO detection, which introduces two new parameters named as initial searching size $K>1$ and deviation factor $\sigma>0$ to characterize the concept of sphere radius, i.e., $D\triangleq\sigma\sqrt{2\ln K}$. Intuitively, compared to the traditional sphere decoding, extra degrees of freedom are obtained in interpreting the decoding trade-off. In particular, by fixing $\sigma$ reasonably, we show that given the sphere radius $D=\sigma\sqrt{2\ln K}$, the number of visited nodes during the searching process is upper bounded by $|S|<nK$, where $n$ is the system dimension. Since ESD actually works the same with Fincke-Pohst SD but with an explicit and tractable decoding trade-off between performance and complexity, the afore-mentioned problem about sphere decoding is addressed.
@@ II-B Fincke-Pohst Sphere Decoding
In this way, the latent interference from elements $x_{i+1},\ldots,x_{n}$ of $\mathbf{x}$ can be alleviated for the decoding of $x_{i}$. Babai’s nearest plane algorithm is also referred to as successive interference cancelation (SIC) algorithm in the field of MIMO detection [6182560].
@@ III-A Algorithm Description
By sharing the same searching space at each layer, the proposed ESD is exactly the same with Fincke-Pohst SD with sphere radius
@@
Although ESD works the same with Fincke-Pohst SD, along the searching process it takes advantages of two parameters $K$ and $\sigma$ rather than the single one $D$, which offers more degrees of freedom to deal with the ILS problem. Typically, it is clear to see that the above equivalence always holds no matter what the deviation factor $\sigma>0$ is, which indicates $\sigma$ should be carefully selected for a better decoding.
@@ III-B Explicit Trade-off Between Performance and Complexity
According to $D=\sigma\sqrt{2\ln K}$ in Theorem 1, extra freedom can be obtained in interpreting the operations of sphere decoding, which provides a feasible way for the analytical diagnose in both decoding performance and complexity.
@@ IV-B Complexity Analysis
Based on Corollary 3 and Theorems 5, when $\sigma=\min_{i}|r_{i,i}|/(2\sqrt{\pi})$, ESD with normalized weighting $p(\cdot)$ achieves a larger shere radius than the one with weighting function $f(\cdot)$ (i.e., Fincke-Pohst SD) under the same complexity upper bound, thus leading to a better decoding trade-off between performance and complexity.
@@ VI-A The Perspective of ESD over Lattice Gaussian Distribution
It is clear that sampling decoding over lattice Gaussian distribution $\mathcal{D}_{\Lambda,\sigma,\mathbf{y}}(\mathbf{x})$ also entails a flexible decoding trade-off determined by the sampling number KK [^2] For straightforward comparison, here we also use $K$ to denote the number of sampling in sampling decoding.. However, compared to the deterministic enumeration in ESD, sampling decoding suffers from inevitable performance loss and complexity waste due to the inherent randomness during the sampling. To make it clear, with $\sigma=\min_{i}|r_{i,i}|/(2\sqrt{\pi})$, the comparison over decoding radius and the number of visited nodes are summarized in Table I. Here, decoding radius is a concept from lattice decoding to evaluate the decoding performance [CongProxity]. Typically, in ESD, decoding radius is the same with sphere radius. As for sampling decoding, ILS or CVP problem will most likely be addressed if $d(\mathbf{\Lambda},\mathbf{y})$ is less than decoding radius while the uncertainty mainly comes from the randomness during the sampling [^3] More details about the decoding radius of sampling decoding can be found in [CongRandom, ZhengWangTIT17].. As can be seen clearly, ESD with $p(\cdot)$ outperforms sampling decoding schemes due to larger decoding radius and less complexity cost. Note that the above decoding radii of ESD algorithms are only based on the survived candidate vectors collected by the pruning threshold, which means the real decoding performance could be better under the help of candidate protection. On the other hand, the computational complexity of randomized sampling decoding is $O(Kn^{2})$, which can serve as an upper bound for the proposed ESD algorithm. Overall, we emphasize that remarkable decoding potential still does exist since ESD with $p(\cdot)$ actually performs the enumeration based on the Gaussian-like distribution $\mathcal{G}(\mathbf{x})$ rather than the Gaussian distribution $\mathcal{D}_{\Lambda,\sigma,\mathbf{y}}(\mathbf{x})$.
@@ VII Simulation
In this section, the performance and the complexity of the proposed ESD algorithm are evaluated by the large-scale MIMO detection. Specifically, given the system model in (2), the $i$th entry of the transmitted signal $\mathbf{x}$, denoted as $x_{i}$, is a modulation symbol taken independently from an $M$-QAM constellation $\mathcal{X}$ with Gray mapping. Meanwhile, we assume a flat fading environment, where the square channel matrix $\mathbf{B}$ contains uncorrelated complex Gaussian fading gains with unit variance and remains constant over each frame duration. Let $E_{b}$ represents the average power per bit at the receiver. Then the signal-to-noise ratio (SNR) $E_{b}/N_{0}=n/(\text{log}_{2}(M)\sigma_{w}^{2})$ where $M$ is the modulation level and $\sigma_{w}^{2}$ is the noise variance. In particular, the ESD and the updated ESD algorithms described in this section are ESD with weighting $f(\cdot)$ and normalized weighting $p(\cdot)$ respectively. Besides, both of them are enhanced by candidate protection, LLL reduction, MMSE augmentation as well as the optimized $\sigma$ through $\alpha_{\text{o}}$ in (91). As a fair comparison, all the other decoding schemes applied here are also strengthened by LLL reduction. Meanwhile, the sampling decoding schemes are also enhanced by MMSE augmentation.
@@
Fig. 7 shows the bit error rate (BER) of the proposed ESD algorithm compared with other decoding schemes in a $12\times 12$ uncoded MIMO system with 64-QAM. Here, lattice-reduction-aided SIC (i.e., Babai’s nearest plane) decoding serves as a performance baseline while ML decoding is implemented by the Schnorr-Euchner (SE) strategy from [DamenDetectionSearch]. Clearly, compared to fixed candidates algorithm (FCA) in [MaxiaoliSoft] and iterative list decoding in [Shimokawa] with 30 samples, sampling decoding algorithms such as Klein’s sampling decoding [Klein], randomized sampling decoding [CongRandom] and IMHK sampling decoding [ZhengWangTIT17] offer not only the improved BER performance but also the promise of smaller sample size $K$. As for the proposed ESD algorithm, it is clear to see that ESD with weighting $f(\cdot)$ is not as good as sampling decoding under the same $K$. This is mainly because the initial searching size $K$ shrinks rapidly by $f(\cdot)$ especially for the limited state space $\mathbf{x}\in\mathcal{X}^{n}$ so that the decoding potential is not well exploited. Nevertheless, a decoding trade-off is still established by ESD with $f(\cdot)$, and one can improve the decoding performance by increasing $K$. As can be seen, there is a remarkable performance gain (i.e., near 2 dB) of ESD with $K=100$ over that with $K=15$. However, since ESD with $f(\cdot)$ is essentially the same with Fincke-Pohst SD, it also implies the decoding trade-off of the conventional sphere decoding is not that charming. On the other hand, as for the updated ESD, i.e., ESD with normalized weighting $p(\cdot)$, substantial performance gain can be found compared to ESD with $f(\cdot)$. Besides, it is clear that the updated ESD outperforms all the sampling decoding schemes under the same size of $K$, which obeys the results shown in Table I. More importantly, the complexity cost of the updated ESD is less than those of sampling decoding schemes, which is illustrated in Fig. 11 and Fig. 12 in detail. Observe that with $K=100$, the performance of the updated ESD suffers negligible loss compared with ML. Therefore, with a moderate $K$, near-ML performance can be achieved.
@@ VIII Conclusions
In this paper, extra degrees of freedom are introduced to sphere decoding for solving the ILS problem in large-scale MIMO detection. Different from the conventional SD, the sphere radius of the proposed ESD algorithm is characterized by the initial searching size $K$ and the deviation factor $\sigma$. Based on it, we showed that the proposed ESD algorithm is exactly the same with the classic Fincke-Pohst SD but with a tractable decoding trade-off between performance and complexity. Moreover, to further exploit the decoding potential, two enhancement mechanisms named as normalized weighting and candidate protection are developed, which not only leads to a better decoding trade-off but also bridges the suboptimal and the optimal decoding performance by simply tuning the initial searching size $K\geq 1$ freely. In addition, further performance enhancement and complexity reduction are investigated to make the proposed ESD algorithm well suited to the various decoding requirements in large-scale MIMO detection.

### v2

@@ Abstract
In this paper, the paradigm of sphere decoding (SD) for solving the integer least square problem (ILS) is revisited, where extra degrees of freedom are introduced to exploit the decoding potential. Firstly, the equivalent sphere decoding (ESD) is proposed, which is essentially the same as the classic Fincke-Pohst sphere decoding but characterizes the sphere radius $D>0$ with two new parameters named as initial searching size $K>1$ and deviation factor $\sigma>0$. By fixing $\sigma$ properly, we show that given the sphere radius $D\triangleq\sigma\sqrt{2\ln K}$, the complexity of ESD in terms of the number of visited nodes is upper bounded by $|S|<nK$, thus resulting in an explicit and tractable decoding trade-off solely controlled by $K$. To the best of our knowledge, this is the first time that the complexity of sphere decoding is exactly specified, where considerable decoding potential can be explored from it. After that, two enhancement mechanisms named as normalized weighting and candidate protection are proposed to further upgrade the ESD algorithm. On one hand, given the same setups of $K$ and $\sigma$, a larger sphere radius is achieved, indicating a better decoding trade-off. On the other hand, the proposed ESD algorithm is generalized, which bridges suboptimal and optimal decoding performance through the flexible choice of $K$. Finally, further performance optimization and complexity reduction with respect to ESD are also derived, and the introduced tractable and flexible decoding trade-off is verified through large-scale MIMO detection.
@@ I Introduction
Nowadays, the large-scale multiple-input multiple-output (MIMO) system has become a promising extension of MIMO in 5G, which boosts the network capacity on a much greater scale without extra bandwidth [5595728, 6736761, 6375940, 6457363]. However, the dramatically increased system size also places a pressing challenge on signal detection, which belongs to integer least square (ILS) problem. As a general way to realize the maximum-likelihood (ML) decoding, the traditional sphere decoding (SD) turns out to be impractical due to the unaffordable complexity in large-scale systems [Kannan, Agrell2002, HassibiExpected, DamenDetectionSearch]. As for those near-ML decoding schemes like fixed-complexity sphere decoding (FCSD), K-best decoder, etc., they are also inapplicable due to the intensive complexity increment and terrible performance deterioration [ZhengWangTSP2019, Hassibisoft, JaldenFSD, 4154776, EmbeddingLuzzi]. To this end, most of related works try to focus on the low-complexity decoding schemes although their performance is severely limited [TabuSrinidhi, DaiL1, GaoX1, MSMIMO1, 6008619, ZhengWangTCOM2019].
@@
First of all, equivalent sphere decoding (ESD) is proposed to solve the ILS problem (i.e., CVP in lattice decoding) in large-scale MIMO detection, which introduces two new parameters named as initial searching size $K>1$ and deviation factor $\sigma>0$ to characterize the concept of sphere radius, i.e., $D\triangleq\sigma\sqrt{2\ln K}$. Intuitively, compared to the traditional sphere decoding, extra degrees of freedom are obtained in interpreting the decoding trade-off. In particular, by fixing $\sigma$ reasonably, we show that given the sphere radius $D=\sigma\sqrt{2\ln K}$, the number of visited nodes during the searching process is upper bounded by $|S|<nK$, where $n$ is the system dimension. Since ESD actually works the same as Fincke-Pohst SD but with an explicit and tractable decoding trade-off between performance and complexity, the afore-mentioned problem about sphere decoding is addressed.
@@ II-B Fincke-Pohst Sphere Decoding
In this way, the latent interference from elements $x_{i+1},\ldots,x_{n}$ of $\mathbf{x}$ can be alleviated for the decoding of $x_{i}$. Babai’s nearest plane algorithm is also referred to as successive interference cancellation (SIC) algorithm in the field of MIMO detection [6182560].
@@ III-A Algorithm Description
By sharing the same searching space at each layer, the proposed ESD is exactly the same as Fincke-Pohst SD with sphere radius
@@
Although ESD works the same as Fincke-Pohst SD, along the searching process it takes advantages of two parameters $K$ and $\sigma$ rather than the single one $D$, which offers more degrees of freedom to deal with the ILS problem. Typically, it is clear to see that the above equivalence always holds no matter what the deviation factor $\sigma>0$ is, which indicates $\sigma$ should be carefully selected for a better decoding.
@@ III-B Explicit Trade-off Between Performance and Complexity
According to $D=\sigma\sqrt{2\ln K}$ in Theorem 1, extra freedom can be obtained in interpreting the operations of sphere decoding, which provides a feasible way for the analytical diagnosis in both decoding performance and complexity.
@@ IV-B Complexity Analysis
Based on Corollary 3 and Theorems 5, when $\sigma=\min_{i}|r_{i,i}|/(2\sqrt{\pi})$, ESD with normalized weighting $p(\cdot)$ achieves a larger sphere radius than the one with weighting function $f(\cdot)$ (i.e., Fincke-Pohst SD) under the same complexity upper bound, thus leading to a better decoding trade-off between performance and complexity.
@@ VI-A The Perspective of ESD over Lattice Gaussian Distribution
It is clear that sampling decoding over lattice Gaussian distribution $\mathcal{D}_{\Lambda,\sigma,\mathbf{y}}(\mathbf{x})$ also entails a flexible decoding trade-off determined by the sampling number KK [^2] For straightforward comparison, here we also use $K$ to denote the number of sampling in sampling decoding.. However, compared to the deterministic enumeration in ESD, sampling decoding suffers from inevitable performance loss and complexity waste due to the inherent randomness during the sampling. To make it clear, with $\sigma=\min_{i}|r_{i,i}|/(2\sqrt{\pi})$, the comparison over decoding radius and the number of visited nodes are summarized in Table I. Here, decoding radius is a concept from lattice decoding to evaluate the decoding performance [CongProxity]. Typically, in ESD, decoding radius is the same as sphere radius. As for sampling decoding, ILS or CVP problem will most likely be addressed if $d(\mathbf{\Lambda},\mathbf{y})$ is less than decoding radius while the uncertainty mainly comes from the randomness during the sampling [^3] More details about the decoding radius of sampling decoding can be found in [CongRandom, ZhengWangTIT17].. As can be seen clearly, ESD with $p(\cdot)$ outperforms sampling decoding schemes due to larger decoding radius and less complexity cost. Note that the above decoding radii of ESD algorithms are only based on the survived candidate vectors collected by the pruning threshold, which means the real decoding performance could be better under the help of candidate protection. On the other hand, the computational complexity of randomized sampling decoding is $O(Kn^{2})$, which can serve as an upper bound for the proposed ESD algorithm. Overall, we emphasize that remarkable decoding potential still does exist since ESD with $p(\cdot)$ actually performs the enumeration based on the Gaussian-like distribution $\mathcal{G}(\mathbf{x})$ rather than the Gaussian distribution $\mathcal{D}_{\Lambda,\sigma,\mathbf{y}}(\mathbf{x})$.
@@ VII Simulation
In this section, the performance and the complexity of the proposed ESD algorithm are evaluated by the large-scale MIMO detection. Specifically, given the system model in (2), the $i$th entry of the transmitted signal $\mathbf{x}$, denoted as $x_{i}$, is a modulation symbol taken independently from an $M$-QAM constellation $\mathcal{X}$ with Gray mapping. Meanwhile, we assume a flat fading environment, where the square channel matrix $\mathbf{B}$ contains uncorrelated complex Gaussian fading gains with unit variance and remains constant over each frame duration. Let $E_{b}$ represent the average power per bit at the receiver. Then the signal-to-noise ratio (SNR) $E_{b}/N_{0}=n/(\text{log}_{2}(M)\sigma_{w}^{2})$ where $M$ is the modulation level and $\sigma_{w}^{2}$ is the noise variance. In particular, the ESD and the updated ESD algorithms described in this section are ESD with weighting $f(\cdot)$ and normalized weighting $p(\cdot)$ respectively. Besides, both of them are enhanced by candidate protection, LLL reduction, MMSE augmentation as well as the optimized $\sigma$ through $\alpha_{\text{o}}$ in (91). As a fair comparison, all the other decoding schemes applied here are also strengthened by LLL reduction. Meanwhile, the sampling decoding schemes are also enhanced by MMSE augmentation.
@@
Fig. 7 shows the bit error rate (BER) of the proposed ESD algorithm compared with other decoding schemes in a $12\times 12$ uncoded MIMO system with 64-QAM. Here, lattice-reduction-aided SIC (i.e., Babai’s nearest plane) decoding serves as a performance baseline while ML decoding is implemented by the Schnorr-Euchner (SE) strategy from [DamenDetectionSearch]. Clearly, compared to fixed candidates algorithm (FCA) in [MaxiaoliSoft] and iterative list decoding in [Shimokawa] with 30 samples, sampling decoding algorithms such as Klein’s sampling decoding [Klein], randomized sampling decoding [CongRandom] and IMHK sampling decoding [ZhengWangTIT17] offer not only the improved BER performance but also the promise of smaller sample size $K$. As for the proposed ESD algorithm, it is clear to see that ESD with weighting $f(\cdot)$ is not as good as sampling decoding under the same $K$. This is mainly because the initial searching size $K$ shrinks rapidly by $f(\cdot)$ especially for the limited state space $\mathbf{x}\in\mathcal{X}^{n}$ so that the decoding potential is not well exploited. Nevertheless, a decoding trade-off is still established by ESD with $f(\cdot)$, and one can improve the decoding performance by increasing $K$. As can be seen, there is a remarkable performance gain (i.e., near 2 dB) of ESD with $K=100$ over that with $K=15$. However, since ESD with $f(\cdot)$ is essentially the same as Fincke-Pohst SD, it also implies the decoding trade-off of the conventional sphere decoding is not that charming. On the other hand, as for the updated ESD, i.e., ESD with normalized weighting $p(\cdot)$, substantial performance gain can be found compared to ESD with $f(\cdot)$. Besides, it is clear that the updated ESD outperforms all the sampling decoding schemes under the same size of $K$, which obeys the results shown in Table I. More importantly, the complexity cost of the updated ESD is less than those of sampling decoding schemes, which is illustrated in Fig. 11 and Fig. 12 in detail. Observe that with $K=100$, the performance of the updated ESD suffers negligible loss compared with ML. Therefore, with a moderate $K$, near-ML performance can be achieved.
@@ VIII Conclusions
In this paper, extra degrees of freedom are introduced to sphere decoding for solving the ILS problem in large-scale MIMO detection. Different from the conventional SD, the sphere radius of the proposed ESD algorithm is characterized by the initial searching size $K$ and the deviation factor $\sigma$. Based on it, we showed that the proposed ESD algorithm is exactly the same as the classic Fincke-Pohst SD but with a tractable decoding trade-off between performance and complexity. Moreover, to further exploit the decoding potential, two enhancement mechanisms named as normalized weighting and candidate protection are developed, which not only leads to a better decoding trade-off but also bridges the suboptimal and the optimal decoding performance by simply tuning the initial searching size $K\geq 1$ freely. In addition, further performance enhancement and complexity reduction are investigated to make the proposed ESD algorithm well suited to the various decoding requirements in large-scale MIMO detection.